One of the excited states of the C molecule has the valence electron configuration 1σg2 2σu*2 1πu1 2πg1. Give the multiplicity and parity of the term.
1
Expert's answer
2015-10-02T03:11:34-0400
Multiplicity is defined by the number of unpaired electrons:
M = 2Sn+1, where S = 1/2, the spin of electron, n – the number of unpaired electrons.
For this configuration, there are 2 unpaired electrons sitting on πu and πg.
Thus, M = 2×0.5×2 + 1 = 3.
The parity of the term is: u×g = u (ungerade). Multiplication u(odd orbital) and g(even orbital) gives u(odd) parity of the term.